# Intrinsic Image Diffusion for Single-view Material Estimation

### 19 Dec 2023 v1

@@ 1 Introduction
Figure 1: Training pipeline. We train a conditional diffusion model to predict albedo and BRDF properties (roughness and metallic) given a single input image. We adapt the learned prior of Stable Diffusion [29] by fine-tuning it on the synthetic InteriorVerse [41] dataset. (i) First, we separately encode the ground-truth (GT) albedo and BRDF properties with a fixed encoder to obtain the material feature maps. We also encode the conditioning image with a trainable encoder. (ii) We add noise to the material features and use our conditional diffusion model to predicted the noise. (iii) The training is supervised with L2 loss between the original and predicted noise. (iv) Using the predicted noise, the predicted material properties can be decoded separately.
@@
    We leverage the real-world image prior of recent diffusion models for material estimation, achieving $77.6\%$ FID and $4.04dB$ PSNR improvement in albedo prediction.
@@ 2 Related work
Material estimation. The field of intrinsic image decomposition in computer graphics and vision has a long and extensive history, with research dating back to the 1970s. Earlier works aimed to decompose an image of a single object [9] into reflectance (albedo) and shading (irradiance) maps using heuristic approaches. Some of these methods include considering the image gradients [16], employing chromaticity clustering [8], utilizing depth cues [5], or incorporating complex priors on the geometry, material, and lighting [3, 2]. [4, 15] took a significant step towards and introduced real-world datasets of indoor scenes for intrinsic image decomposition. These datasets include manually annotated relative comparisons of material properties, providing a valuable benchmark for evaluation. Additionally, [4] proposed a conditional-random-field-based algorithm demonstrating impressive improvement over prior methods.
Recent material estimation methods are usually part of a full inverse rendering framework [42, 6, 36, 25, 24, 1] and often use a learning-based algorithm for complex material predictions trained on increasingly photo-realistic datasets [20, 41] possibly combined with real data [18, 17]. There has been great improvement in the model architectures, including UNet-based [36, 21, 41], or cascaded networks [19, 20], and Transformers [43]. However, all of these methods use a feed-forward model to decompose the input image. In our work, we propose to formulate the task of appearance decomposition probabilistically and use a conditional generative model to sample multiple possible solutions.
A similar line of work focuses on material acquisition, i.e. obtaining high-quality material properties of a single object. Recently, [28] introduced a similar idea for modeling the acquisition from a single scan probabilistically to address the disambiguity between specular and diffuse terms. In turn, our method focuses on the single image multi-object material estimation and utilizes pre-trained image prior to better generalize.
Generative models. The field of generative architectures has recently gained significant attention with diffusion-based models [32, 14, 10, 33, 34, 29] contributing notable advancements to the field as highlighted in the surveys of [38, 26]. These models have demonstrated remarkable fidelity utilizing large-scale text-image datasets [31]. Moreover, these models possess a strong and robust prior that has proven valuable for various downstream applications across different domains [38].
@@ 3.2 Training
Our goal is to utilize their learned prior and adjust it to the material estimation task. We fine-tune the pre-trained text-conditional Stable Diffusion V2 (SD) [29] on our task on synthetic data, similar to [22]. As [29], we follow the latent diffusion architecture with pre-trained frozen encoder $\mathcal{E}$ and decoder $\mathcal{D}$. To condition our diffusion model, we input the image information in two ways to the model. First, we encode it with a trained encoder $\mathcal{E}^{*}$ and concatenate it to the input noisy features. Our encoder $\mathcal{E}^{*}$ has the same architecture as the pre-trained encoder $\mathcal{E}$, but it is initialized randomly. We found that naive downscaling to the latent dimension, as in the original depth-conditioned SD, causes loss in the high-frequency details. Using the frozen pre-trained encoder $\mathcal{E}$ performs worse, suggesting that the task of material estimation requires a different set of features for optimal predictions. Second, we use the CLIP [27, 11] image embedding as a cross-attention conditioner.
@@ 4.1 Synthetic results
Ours - Mean metric 16.14 $\pm$ 3.27 0.73 $\pm$ 0.10 0.26 $\pm$ 0.11
@@
Table 2: Roughness and metallic evaluation. We quantitatively compare the quality of predicted roughness and metallic maps between our method and the baselines [20, 41] on the InteriorVerse dataset [41].
@@
Ours - Mean metric 30.76 $\pm$ 14.05
@@
Diversity. One remaining question about a generative model is whether it produces diverse samples without mode collapse. To evaluate this aspect, we sample $100$ material predictions for a single view (top scene of Fig. 2). We show $4$ predictions in Fig. 3 and visualize the relative variance of the samples. To this end, we normalize the predictions to zero mean and one standard deviation for each channel. Then, we plot the pixel-wise channel-average standard deviation. This variance map can be interpreted as an uncertainty map. We can see that large diffuse objects have stable predictions with a small variance. However, the prediction of the specular, emissive, or small objects is more uncertain, showing complex ambiguity between lighting and material. To quantify this observation for specularity, we investigate the strength of the linear relationship between the albedo variance and metallic properties. We have evaluated the Pearson correlation value per image across all pixels for a subset of the test scenes, which ranges between 0.47 and 0.68, underlining our observation. Note that a perfect correlation should not be expected since specularity is just one source of ambiguity. See qualitative comparison in supplemental.
@@ 4.2 Real-world results
Quantitative evaluation. We report the WHDR metric [4] in Tab. 3, which compares the predicted albedo to human judgments of relative intensities. Our method achieves similar performance quantitatively as the baselines. However, the WHDR metric assesses predictions sparsely and tends to favor smoother outcomes, thus not providing an optimal evaluation of high-frequency details, as shown in Fig. 6 and also found in [37].
@@ 4.3 Ablations
Figure 8: Model ablation. Leveraging Stable Diffusion’s [29] learned prior of real images significantly improves our material diffusion.
@@ D.2 Real Results
Image reconstruction We thank the authors of [41] for providing the code for running their method and for discussing the results.

### 21 Mar 2024 v2

# Intrinsic Image Diffusion for Indoor Single-view Material Estimation

@@ 1 Introduction
Figure 1: Training pipeline. We train a conditional diffusion model to predict albedo and BRDF properties (roughness and metallic) given a single input image. We adapt the learned prior of Stable Diffusion [36] by fine-tuning it on the synthetic InteriorVerse [48] dataset. Models being trained are marked with yellow. (i) First, we separately encode the ground-truth (GT) albedo and BRDF properties with a fixed encoder to obtain the material feature maps. We also encode the conditioning image with a trainable encoder. (ii) We add noise to the material features and use our conditional diffusion model to predicted the noise. (iii) The training is supervised with L2 loss between the original and predicted noise. (iv) Using the predicted noise, the predicted material properties can be decoded separately.
@@
    We leverage the real-world image prior of diffusion models for material estimation, achieving $77.6\%$ FID and $4.04dB$ PSNR improvement in albedo prediction.
  3. 3.
    Using our material predictions, we optimize for lighting as multiple point lights and a global environment map.
@@ 2 Related work
Material estimation. The field of intrinsic image decomposition in computer graphics and vision has a long and extensive history, with research dating back to the 1970s. Earlier works aimed to decompose an image of a single object [14] into reflectance (albedo) and shading (irradiance) maps using heuristic approaches. Some of these methods include considering the image gradients [22, 18], employing chromaticity clustering [13], utilizing depth cues [8], or incorporating complex priors on the geometry, material, and lighting [3, 2]. [4, 21] took a significant step towards and introduced real-world datasets of indoor scenes for intrinsic image decomposition. These datasets include manually annotated relative comparisons of material properties, providing a valuable benchmark for evaluation. Additionally, [4] proposed a conditional-random-field-based algorithm demonstrating impressive improvement over prior methods.
Recent material estimation methods are usually part of a full inverse rendering framework [49, 9, 43, 32, 31, 1] and often use a learning-based algorithm for complex material predictions trained on increasingly photo-realistic datasets [27, 48] possibly combined with real data [25, 24]. There has been great improvement in the model architectures, including UNet-based [43, 28, 48], or cascaded networks [26, 27], and Transformers [50]. However, all of these methods use a feed-forward model to decompose the input image. We propose to formulate the task of appearance decomposition probabilistically and use a conditional generative model to sample multiple possible solutions.
Another line of work focuses on material acquisition, i.e. obtaining high-quality material properties of a single object. [35] introduced a similar idea for modeling the acquisition from a single scan probabilistically to address the disambiguity between specular and diffuse terms. Our method focuses on the single image multi-object material estimation and utilizes pre-trained image prior to better generalize.
Generative models. The field of generative architectures has recently gained significant attention with diffusion-based models [39, 20, 15, 40, 41, 36] contributing notable advancements to the field as highlighted in the surveys of [45, 33]. These models have demonstrated remarkable fidelity utilizing large-scale text-image datasets [38]. Moreover, these models possess a strong and robust prior that has proven valuable for various downstream applications across different domains [45], such as relighting [5] and even for extracting intrinsic properties [10, 23, 7] determinisically.
@@ 3.2 Training
Our goal is to utilize their learned prior and adjust it to the material estimation task. We fine-tune the pre-trained text-conditional Stable Diffusion V2 (SD) [36] on our task on synthetic data, similar to [29]. As [36], we follow the latent diffusion architecture with pre-trained frozen encoder $\mathcal{E}$ and decoder $\mathcal{D}$. To condition our diffusion model, we input the image information in two ways to the model. First, we encode it with a trained encoder $\mathcal{E}^{*}$ and concatenate it to the input noisy features. Our encoder $\mathcal{E}^{*}$ has the same architecture as the pre-trained encoder $\mathcal{E}$, but it is initialized randomly and gives $3$-channel features; thus, the overall input channel size is $11$. We found that naive downscaling to the latent dimension, as in the original depth-conditioned SD, causes loss in the high-frequency details. Using the frozen pre-trained encoder $\mathcal{E}$ performs worse, suggesting that the task of material estimation requires a different set of features for optimal predictions. Second, we use the CLIP [34, 16] image embedding as a cross-attention conditioner.
@@ 4.1 Synthetic results
Ours - Mean metric 16.14 $\pm$ 3.27 0.73 $\pm$ 0.10 0.26 $\pm$ 0.11 |
@@
Table 2: Roughness and metallic evaluation. We compare the quality of predicted roughness and metallic maps between our method and the baselines [27, 48] on the InteriorVerse dataset [48].
@@
Ours - Mean metric 30.76 $\pm$ 14.05 |
@@
Diversity. One remaining question about a generative model is whether it produces diverse samples without mode collapse. To evaluate this aspect, we sample $100$ material predictions for a single view (top scene of Fig. 2). We show $4$ predictions in Fig. 3 and visualize the relative variance of the samples. To this end, we normalize the predictions to zero mean and one standard deviation for each channel. Then, we plot the pixel-wise channel-average standard deviation. This variance map can be interpreted as an uncertainty map. We can see that large diffuse objects have stable predictions with a small variance. However, the prediction of the specular, emissive, or small objects is more uncertain, showing complex ambiguity between lighting and material. To quantify this observation for specularity, we investigate the strength of the linear relationship between the albedo variance and metallic properties. We have evaluated the Pearson correlation value per image across all pixels for a subset of the test scenes, which ranges between 0.47 and 0.68, underlining our observation. Note that perfect correlation should not be expected as specularity is just one source of ambiguity. See qualitative comparison in supplemental.
@@ 4.2 Real-world results
Quantitative evaluation. We report the WHDR metric [4] in Tab. 3, which compares the predicted albedo to human judgments of relative intensities. Our method achieves similar performance quantitatively as the baselines. However, the WHDR metric assesses predictions sparsely and tends to favor smoother outcomes, thus not providing an optimal evaluation of high-frequency details, as shown in Fig. 6 and also found in [44, 6, 12].
@@ 4.4 Limitations and Future Work
Figure 8: Model ablation. Leveraging Stable Diffusion’s [36] real image prior significantly improves our material diffusion.
@@ D.2 Real Results
Image reconstruction We provide additional image rerenderings using our full pipeline in Fig. 11. We thank the authors of [48] for providing the code for running their method and for discussing the results.
S-AWARE Network [18] We compare against S-AWARE [18] in Fig. 12. Our method predicts roughness and metallic properties as well and improves upon the albedo estimation by avoiding baked-in lighting or shadows. We thank the authors for providing us with their results.
@@
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Image reconstruction. Additional image reconstruction results using our full pipeline.
@@ Appendix E Additional Ablations
Input
S-AWARE [18]
Ours
Figure 12: Comparison to S-AWARE [18]. Our method gives complex material maps and also improved albedo without baked-in lighting and shadows.